OpenStudy (anonymous):

Let f(x)= square root of x-3 and g(x)=x^2+5. (f+g)(-3). Thanks for the help :)

jimthompson5910 (jim_thompson5910):

plug in x = -3 into each function then evaluate finally, add up the two results

OpenStudy (anonymous):

I got (square root of 6+ 14)

jimthompson5910 (jim_thompson5910):

is the function f(x) this \[\large f(x) = \sqrt{x-3}\] OR is it this \[\large f(x) = \sqrt{x}-3\]

OpenStudy (anonymous):

the top one

jimthompson5910 (jim_thompson5910):

\[\large f(x) = \sqrt{x-3}\] \[\large f(-3) = \sqrt{-3-3}\] \[\large f(-3) = \sqrt{-6}\]

jimthompson5910 (jim_thompson5910):

at this point, have you learned about complex or imaginary numbers?

OpenStudy (anonymous):

Yeah I got that

OpenStudy (anonymous):

Yes I have

jimthompson5910 (jim_thompson5910):

ok so \[\large f(x) = \sqrt{x-3}\] \[\large f(-3) = \sqrt{-3-3}\] \[\large f(-3) = \sqrt{-6}\] \[\large f(-3) = i*\sqrt{6}\]

jimthompson5910 (jim_thompson5910):

g(-3)=(-3)^2 + 5 = 9+5 = 14 g(-3) = 14 --------------------------- (f+g)(-3) = f(-3) + g(-3) \[\large (f+g)(-3) = i\sqrt{6} + 14\] \[\large (f+g)(-3) = 14 + i\sqrt{6}\]

OpenStudy (anonymous):

Thank you

jimthompson5910 (jim_thompson5910):

yw